# Yi Lin 11 $0$-0001-8443-2246    Lucas M. Tabajara Currently at Runtime Verification, Inc.11 $0$-0001-9608-1404    Moshe Y. Vardi 11 $0$-0002-0661-5773

### v2

@@ 2.2 Dynamic Programming Concepts - Project-Join Trees
(a) A Graded Project-Join Tree Instance, $e$ denotes labels
(b) Intermediate tree 1
(c) Intermediate tree 2
(d) Intermediate tree 3
Figure 1: Original and Intermediate Trees Generated in Algorithms for the CNF Example $(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(\neg x_{1}\lor x_{2}\lor y_{6})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\land(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})$
3 Realizability Checking [^2] Proofs for all lemmas and theorems can be found in the Appendix 0.A.
@@ 3.1 Theoretical Basis and Valuations in Trees
As for the CNF with the graded project-join tree in Figure 1(a), the pre and post valuations for nodes $1,4,5,2,3$ are equivalent to the BDDs representing their original clauses. $\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=\llbracket(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\rrbracket,\texttt{BV}_{\texttt{post}}(\mathcal{T},6)=(\exists y_{4})\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=(x_{1}\lor\neg y_{5}\lor\neg x_{3}),\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(x_{1}\lor\neg y_{5}\lor\neg x_{3}),\texttt{BV}_{\texttt{post}}(\mathcal{T},7)=(\exists y_{5})\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=\llbracket 1\rrbracket,\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=(\neg x_{1}\lor x_{2}\lor y_{6}),\texttt{BV}_{\texttt{post}}(\mathcal{T},9)=(\exists y_{6})\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=\llbracket 1\rrbracket$.
@@ 3.2 Determining Nullary, Partial and Full Realizability
In the case of the example formula in Figure 1(a), once we obtain the pre and post valuations from Algorithm 1, $B_{pureX}=\llbracket 1\rrbracket$ and $b=1$ from the post valuations $\texttt{BV}_{\texttt{post}}(\mathcal{T},7)$ and $\texttt{BV}_{\texttt{post}}(\mathcal{T},9)$ on XLeaves. Hence we get full realizability.
@@ 19             $b\leftarrow 0$
To illustrate by the example in Figure 1(a), for this formula, we do not check in Algorithm 3 because by Algorithm 2 full realizability is returned.
@@ 4.1 Monolithic Approach
The above monolithic synthesis framework was generalized in [27] to the case where the formula $\varphi$ is given as a conjunction of factors. A factor can be a formula, for example, a single clause or a conjunction of clauses (called a “cluster” in that work). In that case, the principle of early quantification mentioned in Section 3.1 can be applied. See Appendix 0.B for details. As shown in [27], the tool, Factored RSynth, generally outperforms RSynth.
@@ 4.2 Synthesis Using Graded Project-Join Trees
We still give a running example for the problem in Figure 1(a). Once full realizability is detected, we apply Algorithm 4 and construct the witnesses in the top-down manner. First, we get the witness $g_{6}=1$ for $y_{6}$ by $\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=\llbracket(\neg x_{1}\lor x_{2}\lor y_{6})\rrbracket$. Then construct the witness $g_{5}=(x_{1}\land x_{2})\land\neg x_{3}$ for $y_{5}$ by substituting $y_{5}=1$ in $\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=\llbracket(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(x_{1}\lor\neg y_{5}\lor\neg x_{3})\rrbracket$. After that we go to node $6$ and get $g_{4}=x_{1}\lor\neg x_{3}$ from $\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=\llbracket(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\rrbracket$. Then since for this specific CNF $g_{5}$ is not dependent on $g_{4}$, we do not have additional substitutions. The witnesses are correct by $\varphi[y_{4}\mapsto g_{4}][y_{5}\mapsto g_{5}][y_{6}\mapsto g_{6}]=1$.
@@ 6 Concluding Remarks
To summarize the contribution in this work, we propose a novel symbolic dynamic programming approach for realizability checking and witness construction in boolean synthesis, based on graded project-join trees. The algorithm we propose here combine a bottom-up realizability-checking phase with a top-down synthesis phase. We demonstrated experimentally that our approach, implemented in the DPSynth tool, is powerful and more scalable than the approach based on CSP heuristics (Factored RSynth). Another crucial contribution of this work is the inclusion of partial realizability checking, which applies to 30% of the total number of benchmarks. As we explain in introduction, this consideration, which is lacking in other works, is motivated by the need in modular circuit design and temporal synthesis to locate the scope of realizable inputs in iterative constructions.

### v3

@@ 2.2 Dynamic Programming Concepts - Project-Join Trees
Figure 1 shows an example Graded Project-Join Tree for a CNF formula, along with intermediate trees produced in the course of the execution of our synthesis algorithm. In the upcoming sections we will refer back to this example to illustrate the individual steps of the algorithm related to each intermediate tree.
(a) A Graded Project-Join Tree Instance
(b) Tree after evaluating node 6
(c) Tree after evaluating node 7
(d) Tree after evaluating node 9
Figure 1: Original and intermediate trees generated by our algorithms for the CNF example $(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(\neg x_{1}\lor x_{2}\lor y_{6})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\land(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})$. The label for each internal node is denoted by $e$.
3 Realizability Checking [^2] Proofs for all lemmas and theorems can be found in the Appendix A.
@@ 3.1 Theoretical Basis and Valuations in Trees
For the CNF with the graded project-join tree in Figure 1(a), the pre- and post-valuations for nodes $1,4,5,2,3$ are equivalent to the BDDs representing their original clauses. $\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=\llbracket(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\rrbracket,\texttt{BV}_{\texttt{post}}(\mathcal{T},6)=(\exists y_{4})\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=(x_{1}\lor\neg y_{5}\lor\neg x_{3}),\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(x_{1}\lor\neg y_{5}\lor\neg x_{3}),\texttt{BV}_{\texttt{post}}(\mathcal{T},7)=(\exists y_{5})\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=\llbracket 1\rrbracket,\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=(\neg x_{1}\lor x_{2}\lor y_{6}),and\texttt{BV}_{\texttt{post}}(\mathcal{T},9)=(\exists y_{6})\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=\llbracket 1\rrbracket$.
@@ 3.2 Determining Nullary, Partial and Full Realizability
In the case of the example formula in Figure 1, once we obtain the pre- and post-valuations from Algorithm 1, $B_{pureX}=\llbracket 1\rrbracket$ and $b=1$ from the post-valuations $\texttt{BV}_{\texttt{post}}(\mathcal{T},7)$ (Figure 1(c)) and $\texttt{BV}_{\texttt{post}}(\mathcal{T},9)$ (Figure 1(d)) on XLeaves. Hence we get full realizability.
@@ 19             $b\leftarrow 0$
In the case of the example in Figure 1, for this formula, we do not need to check Algorithm 3 because by Algorithm 2 full realizability is returned.
@@ 4.1 Monolithic Approach
The above monolithic synthesis framework was generalized in [27] to the case where the formula $\varphi$ is given as a conjunction of factors. A factor can be a formula, for example, a single clause or a conjunction of clauses (called a “cluster” in that work). In that case, the principle of early quantification mentioned in Section 3.1 can be applied. See Appendix B for details. As shown in [27], the tool, Factored RSynth, generally outperforms RSynth.
@@ 4.2 Synthesis Using Graded Project-Join Trees
Continuing with the running example for the problem in Figure 1, once full realizability is detected, we apply Algorithm 4 and construct the witnesses in top-down manner. First, we get the witness $g_{6}=1$ for $y_{6}$ by $\texttt{BV}_{\texttt{pre}}(\mathcal{T},9)=\llbracket(\neg x_{1}\lor x_{2}\lor y_{6})\rrbracket$. Then we construct the witness $g_{5}=(x_{1}\land x_{2})\lor\neg x_{3}$ for $y_{5}$ by substituting $y_{5}=1$ in $\texttt{BV}_{\texttt{pre}}(\mathcal{T},7)=\llbracket(x_{1}\lor\neg x_{2}\lor x_{3}\lor y_{5})\land(\neg x_{3}\lor x_{2}\lor\neg y_{5})\land(x_{1}\lor\neg y_{5}\lor\neg x_{3})\rrbracket$. After that, we go to node $6$ and get $g_{4}=x_{1}\lor\neg x_{3}$ from $\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)=\llbracket(x_{1}\lor y_{4}\lor\neg y_{5})\land(\neg x_{3}\lor x_{1}\lor\neg y_{4})\rrbracket$. By the algorithm, we would need to substitute $g_{5}$ into $\texttt{BV}_{\texttt{pre}}(\mathcal{T},6)$ before computing $g_{4}$, but since for this specific CNF $g_{4}$ is not actually dependent on $g_{5}$, this does not change the witness. The witnesses are correct by $\varphi[y_{4}\mapsto g_{4}][y_{5}\mapsto g_{5}][y_{6}\mapsto g_{6}]=1$.
@@ 6 Concluding Remarks
To summarize the contribution in this work, we propose a novel symbolic dynamic programming approach for realizability checking and witness construction in boolean synthesis, based on graded project-join trees. The algorithm we propose here combine a bottom-up realizability-checking phase with a top-down synthesis phase. We demonstrated experimentally that our approach, implemented in the DPSynth tool, is powerful and more scalable than the approach based on CSP heuristics (Factored RSynth). Another crucial contribution of this work is the inclusion of partial realizability checking, which applies to 30% of the total number of benchmarks. As we explain in introduction, this consideration is motivated by the need in modular circuit design and temporal synthesis to locate the scope of realizable inputs in iterative constructions.